OpenStudy (anonymous):

Assuming x > 0, which of these expressions is equivalent to 11 times the square root of 245 x to the third plus 9 times the square root of 45 x to the third? ?

OpenStudy (jack1):

heya

OpenStudy (anonymous):

Hey :)

OpenStudy (jack1):

...which of which expressions dude, is there more to the q...?

OpenStudy (anonymous):

Umm...let me see.. 5 x times the square root of 104 x 20 times the square root of 290 x to the sixth 20 x times the square root of 290 x 104 x times the square root of 5 x

OpenStudy (jack1):

k so use BODMAS rules to kinda guess the order of operations here: i think the question looks something like:

OpenStudy (jack1):

11 times the square root of 245 x to the third plus 9 times the square root of 45 x to the third? ? \[11 \times \sqrt{245x ^{3}} + \sqrt{45x ^{3}}\]

OpenStudy (jack1):

but it may be ... (245x)^3 and (45x)^3 ... which changes everything but lets assume the first guess is correct and see how we go

OpenStudy (anonymous):

Here ill type it up. :)

OpenStudy (jack1):

cool

OpenStudy (anonymous):

\[11\sqrt{245^3} + 9\sqrt{45^3}\] There ya go^^

OpenStudy (jack1):

ah, nice, i missed the 9, cheers man

OpenStudy (anonymous):

lol your helping me a ton already thanks ^-^

OpenStudy (jack1):

so, starting at the back, the factors of 45 are: 9 and 5 so sqrt 45 = (sqrt 5) times (sqrt 9) = (sqrt 5) times (3) \[=3 \sqrt{5}\]

OpenStudy (jack1):

now sqrt is the same as (to the power of half) so \[\sqrt{x ^{3}} = (x ^{3})^{0.5} = (x ^{3})^{1/2} = x ^{(3\times(1/2)}\] \[= x ^{3/2}\]

OpenStudy (jack1):

so expanding that out youve got 9 * 3 * x^(3/2) * sqrt 5 = 27 * sqrt 5 * x^(3/2)

OpenStudy (anonymous):

You may have just lost me lol :P

OpenStudy (jack1):

bear with me, it'll come together and make sense at the end, promise

OpenStudy (anonymous):

lol k :D

OpenStudy (jack1):

now the start factors of 245 = 5 and 49 (5*49=245) and 49 = 7 *7 = 7^2 so sqrt of 245 = sqrt 49 times sqrt 5 = 7 times sqrt 5 \[=7\sqrt{5}\]

OpenStudy (jack1):

so 11 times sqrt (245x^3) = 11 * 7 * sqrt 5 * x^(3/2)

OpenStudy (jack1):

\[x ^{(3/2)} = x \times \sqrt{x}\]

OpenStudy (jack1):

so bring it all together now:

OpenStudy (anonymous):

20080√5x^3/2??

OpenStudy (jack1):

11*7 = 77 so 77x * sqrt (5x) plus 3*9 = 27 so 27x *sqrt (5x) so final answer should be: 77x * sqrt (5x) + 27x *sqrt (5x)

OpenStudy (anonymous):

That ones not up there on the list :'(

OpenStudy (jack1):

if you add the like terms tho it becomes 104x * sqrt (5x)

OpenStudy (jack1):

you can add the 27 and the 77 together as they are both to the same x value

OpenStudy (jack1):

\[104x \times \sqrt (5x) \] is one of your answers tho... ;D

OpenStudy (anonymous):

Ohhh I see NOW!!!

OpenStudy (jack1):

sweet! all good, need any clarification on any of the steps dude...?

OpenStudy (anonymous):

Nah bro I got it when I look back at it now! :)

OpenStudy (jack1):

noice, slaters then!